solve : 4x2+2x+3=0

Grade:12

1 Answers

Rajendra Sharma
11 years ago

x^2 + (1/2)x = -(3/4)

(x+1/4)^2 = 1/16 - 3/4 = -11/16 implies x = (-1+i(11)^(1/2))/4 and its conjugate.